# Einstein metrics on homogeneous superspaces

### v1

@@ 3.4 Proof of the Ricci curvature formula
$\quad+2\sum_{j=1}^{d}\Big{(}\langle[[\widebar{X}_{j},X_{i}]_{\mathfrak{k}},\widebar{X}_{i}]_{\mathfrak{m}},X_{j}\rangle+\langle[[X_{i},\widebar{X}_{i}]_{\mathfrak{m}},\widebar{X}_{j}]_{\mathfrak{m}},X_{j}\rangle-\langle[[X_{i},\widebar{X}_{j}]_{\mathfrak{k}},X_{j}]_{\mathfrak{m}},\widebar{X}_{i}\rangle$
$\qquad\qquad\quad-\langle[X_{i},[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{m}}]_{\mathfrak{m}},X_{j}\rangle-\langle X_{i},[[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{k}},X_{j}]_{\mathfrak{m}}\rangle+\langle[[\widebar{X}_{j},X_{i}]_{\mathfrak{g}},\widebar{X}_{i}]_{\mathfrak{m}},X_{j}\rangle\Big{)}.$ (3.22)
Apply the super-Jacobi identity to the last expression, and use (3.16) to conclude that
$\sum_{j=1}^{d}\,\langle[[X_{i},\widebar{X}_{i}]_{\mathfrak{g}},\widebar{X}_{j}]_{\mathfrak{m}},X_{j}\rangle=\sum_{j=1}^{d}\,\langle[[X_{i},\widebar{X}_{i}]_{\mathfrak{m}},\widebar{X}_{j}]_{\mathfrak{m}},X_{j}\rangle.$
We can then rewrite the $j$-sum with the six terms in (3.4) as
$2\sum_{j=1}^{d}\Big{(}\langle[[\widebar{X}_{j},X_{i}]_{\mathfrak{k}},\widebar{X}_{i}]_{\mathfrak{m}},X_{j}\rangle-\langle[[X_{i},\widebar{X}_{j}]_{\mathfrak{k}},X_{j}]_{\mathfrak{m}},\widebar{X}_{i}\rangle+\langle[X_{i},[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{k}}]_{\mathfrak{m}},X_{j}\rangle-\langle X_{i},[[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{k}},X_{j}]_{\mathfrak{m}}\rangle\Big{)},$
which is zero due to $\operatorname{ad}_{\mathfrak{k}}$-invariance. Since $B(X_{i},\widebar{X}_{i})=(-1)^{[i]}B(\widebar{X}_{i},X_{i})$, the formula in Theorem 3.11 follows. ∎
@@ 3.5 Naturally reductive metrics
We say that the $G$-invariant metric $\langle$-,$\rangle$ - , - ⟩ on the homogeneous superspace $M=G/K$ is naturally reductive if
@@
for all $X,Y,Z\in\mathfrak{m}$. This mimics the definition of a naturally reductive metric in classical homogeneous geometry [DZ79]. Such metrics have been used extensively in the study of the Einstein equation, spectral geometry, the prescribed Ricci curvature problem, and other topics; see, e.g., [DZ79, GS10, Lau20, AGP20]. When $\langle$-,$\rangle$ - , - ⟩ is naturally reductive, the curvature formulas obtained above simplify substantially.
@@
Let $\langle$-,$\rangle$ - , - ⟩ be naturally reductive. Then, $U=0$.
@@
Since this holds for all $Z$, the result follows from the non-degeneracy of $\langle$-,$\rangle$ - , - ⟩. ∎
@@
Let $M=G/K$ be a homogeneous superspace with $G$-invariant graded naturally reductive Riemannian metric $\langle$-,$\rangle$ - , - ⟩, where $G=(G_{0},\mathfrak{g})$ is a connected Lie supergroup, $\mathfrak{g}$ a compact real form of a basic classical Lie superalgebra, and $K$ a connected closed Lie subsupergroup of $G$. Let $\{X_{1},\ldots,X_{d}\}$ be a $g$-normalised homogeneous basis for $\mathfrak{m}=T_{K}(M)$, with right dual basis $\{\widebar{X}_{1},\ldots,\widebar{X}_{d}\}$. Then, for every $i,j,k,\ell\in\{1,\ldots,d\}$,
@@ 3.6 Diagonal metrics
Recall that $\mathfrak{g}$ admits the $Q$-orthogonal decomposition $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{m}$. As $K$ is connected, this means that $\mathfrak{m}$ is an $\operatorname{ad}_{\mathfrak{k}}$-representation, that is, $[\mathfrak{k},\mathfrak{m}]\subseteq\mathfrak{m}$. In the following, we will assume $\mathfrak{m}$ decomposes into $Q$-orthogonal $\operatorname{ad}_{\mathfrak{k}}$-irreducible representations,
@@
As we will discuss in the following, the curvature formulas obtained in Section 3.3 simplify quite dramatically for diagonal metrics. We will frequently use an ordered basis for $\mathfrak{m}$ that is adapted to the decomposition (3.24). For each $i=1,\ldots,s$, we thus let $\mathfrak{M}_{i}=\{e^{i}_{\alpha}\mid\alpha=1,\ldots,\dim(\mathfrak{m}_{i})\}$ be an ordered $Q$-normalised homogeneous basis for $\mathfrak{m}_{i}$ such that the corresponding matrix representation of $Q$ is of the form (3.11), and let $\{\bar{e}_{\alpha}^{i}\mid\alpha=1,\ldots,\dim(\mathfrak{m}_{i})\}$ be the right $Q$-dual basis, defined by
@@
Let $g$ be a $G$-invariant graded Riemannian metric on $M$. Then, by Theorem 3.3, $g$ is identified with an $\operatorname{Ad}_{K}$-invariant scalar superproduct $\langle$-,$\rangle$ - , - ⟩ on $\mathfrak{m}$. Since $K$ is connected, the $\operatorname{Ad}_{K}$-invariance is equivalent to the $\operatorname{ad}_{\mathfrak{k}}$-invariance of $\langle$-,$\rangle$ - , - ⟩ at the level of Lie superalgebras. By Schur’s lemma, the metric $g$ is therefore of the form
$g=\langle-,-\rangle=\sum_{i=1}^{s}x_{i}Q|_{\mathfrak{m}_{i}},\qquad x_{1},\ldots,x_{s}\in\mathbb{R}^{\times},$ (3.26)
@@ 3.6.1 Structure constants
Using $[\bar{e}^{j}_{\beta},\bar{e}_{\alpha}^{i}]_{\mathfrak{m}_{k}}=\sum_{\gamma\in I^{k}}Q([\bar{e}^{j}_{\beta},\bar{e}_{\alpha}^{i}]_{\mathfrak{m}},\bar{e}_{\gamma}^{k})e_{\gamma}^{k}$, we see that
@@
It remains to show that $[ijk]$ is independent of the choice of basis for $\mathfrak{m}_{i}$. Suppose there is another $Q$-normalised homogeneous basis $\{f_{\beta}^{i}\mid\beta\in I^{i}\}$ for $\mathfrak{m}_{i}$, with right dual basis $\{\bar{f}_{\beta}^{i}\mid\beta\in I^{i}\}$, and let $A^{i}$ and $B^{i}$ be the (even) transition matrices such that
@@
Since $\mathfrak{g}$ admits the $Q$-orthogonal decomposition $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{m}$ and $[\mathfrak{k},\mathfrak{m}]\subseteq\mathfrak{m}$, we have
@@ 3.6.2 Ricci curvature
For diagonal metrics, we now derive expressions for the Ricci coefficients $r_{i}$, $i=1,\ldots,s$, in (3.27). Although the superdimensions $d_{i}$ may be 00 (unlike the corresponding ordinary dimensions), the formula for the Ricci coefficients in Theorem 3.24 below is analogous to the corresponding classical result; cf. [PS97, Lemma 1.1] and [PR19, Lemma 3.3]. If $\dim(\mathfrak{m}_{\bar{1}})=0$, then Theorem 3.24 gives a Ricci curvature formula for a pseudo-Riemannian manifold.
@@
For each $i=1,\ldots,s$, let $\{X^{i}_{\alpha}\mid\alpha\in I^{i}\}$ be a $g$-normalised basis for $\mathfrak{m}_{i}$, with right $g$-dual basis $\{\widebar{X}_{\alpha}^{i}\mid\alpha\in I^{i}\}$, and let $\{X_{1},\dots,X_{d}\}$ be a corresponding basis for $\mathfrak{m}$ adapted to the decomposition (3.24). Then, $Q$-orthogonality implies that $\langle X^{i}_{\alpha},X^{j}_{\beta}\rangle=0$ for $i\neq j$ and any $\alpha,\beta$. For each $i=1,\ldots,s$, we also introduce the renormalised basis $\{e^{i}_{\alpha}=\sqrt{\tilde{x}_{i}}X_{\alpha}^{i}\mid\alpha\in I^{i}\}$ for $\mathfrak{m}_{i}$, whose right $Q$-dual basis is seen to be given by $\{\bar{e}_{\alpha}^{i}=\epsilon_{i}\sqrt{\tilde{x}_{i}}\widebar{X}^{i}_{\alpha}\mid\alpha\in I^{i}\}$.
@@
where the second, second-to-last and last equalities follow from $[X_{\alpha}^{i},X_{j}]_{\mathfrak{m}}=\sum_{k=1}^{d}\,\langle X_{k},[X_{\alpha}^{i},X_{j}]_{\mathfrak{m}}\rangle\,\widebar{X}_{k}$, $[\bar{e}_{\beta}^{j},\bar{e}^{i}_{\alpha}]_{\mathfrak{m}_{k}}=\sum_{\gamma\in I^{k}}Q\big{(}[\bar{e}_{\beta}^{j},\bar{e}^{i}_{\alpha}]_{\mathfrak{m}_{k}},\bar{e}_{\gamma}^{k}\big{)}e_{\gamma}^{k}$ and (3.6.1), respectively. Taking the graded sum over the $Q$-normalised homogeneous basis for $\mathfrak{m}_{i}$, we obtain
@@
In the homogeneous $\mathfrak{m}_{i}$-basis $\{e^{i}_{\alpha}\mid\alpha\in I^{i}\}$ introduced in the proof of Theorem 3.24, there exists at least one even vector, because $d_{i}=0$. Let us select one such vector and denote it by $e_{\alpha}^{i}$. Since (3.35) holds, we have
@@ 4 Flag supermanifolds
In this section, we introduce a class of homogeneous superspaces called flag supermanifolds. In analogy with the construction of generalised flag manifolds in terms of so-called painted Dynkin diagrams in the non-super case [Arv03], we introduce what we call circled Dynkin diagrams to describe the flag supermanifolds in Section 4.1. In Section 4.2, we identify the compact real forms of the Lie superalgebras of types $A$ and $C$. In Section 4.3, Section 4.4 and Section 4.5, we then classify Einstein metrics on flag supermanifolds constructed from circled Dynkin diagrams of types $A$ and $C$.
@@ 4.1 Circled Dynkin diagrams
$\mathfrak{g}_{\alpha}^{\mathbb{C}}:=\big{\{}X\in\mathfrak{g}^{\mathbb{C}}\mid[H,X]=\alpha(H)X,\ \text{for all}\ H\in\mathfrak{h}^{\mathbb{C}}\big{\}}.$
@@
$\mathfrak{m}=\operatorname{span}_{\mathbb{R}}\big{\{}X\in\mathfrak{m}_{\bar{0}}^{\mathbb{C}}\mid X^{*}=-X\big{\}}\oplus\operatorname{span}_{\mathbb{R}}\big{\{}\sqrt{\imath}X\in\mathfrak{m}_{\bar{1}}^{\mathbb{C}}\mid X^{*}=-X\big{\}},$ (4.3)
@@
$\{A_{\alpha},B_{\alpha}\mid\alpha\in\Delta_{M}\cap\Delta_{\bar{0}}^{+}\}\cup\{\sqrt{\imath}A_{\alpha},\sqrt{\imath}B_{\alpha}\mid\alpha\in\Delta_{M}\cap\Delta_{\bar{1}}^{+}\}.$
In the non-super case, the above construction is encoded in a so-called painted Dynkin diagram, where the nodes of $\Pi_{M}$ are painted black while the other nodes of $\Pi$ remain white. In the super-setting, however, black nodes typically represent non-isotropic odd roots, so to avoid confusion, we shall circle the nodes of $\Pi_{M}$, thereby obtaining a circled Dynkin diagram of $G$. The uncircled nodes of $\Pi$ form a simple root system of $\mathfrak{k}_{\mathrm{ss}}^{\mathbb{C}}$.

### v2

@@ 3.4 Proof of the Ricci curvature formula
$+2\sum_{j=1}^{d}\Big(\langle[[\widebar{X}_{j},X_{i}]_{\mathfrak{g}},\widebar{X}_{i}]_{\mathfrak{m}},X_{j}\rangle+\langle[[X_{i},\widebar{X}_{i}]_{\mathfrak{m}},\widebar{X}_{j}]_{\mathfrak{m}},X_{j}\rangle$
$\qquad\qquad-\langle[X_{i},[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{m}}]_{\mathfrak{m}},X_{j}\rangle-\langle X_{i},[[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{k}},X_{j}]_{\mathfrak{m}}\rangle\Big).$ (3.22)
We can use the super-Jacobi identity
$[[\widebar{X}_{j},X_{i}]_{\mathfrak{g}},\widebar{X}_{i}]_{\mathfrak{m}}=-[[X_{i},\widebar{X}_{i}]_{\mathfrak{g}},\widebar{X}_{j}]_{\mathfrak{m}}+[X_{i},[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{g}}]_{\mathfrak{m}}$
to rewrite the first term in the last $j$-sum, and use (3.16) to conclude that
$\sum_{j=1}^{d}\,\langle[[X_{i},\widebar{X}_{i}]_{\mathfrak{g}},\widebar{X}_{j}]_{\mathfrak{m}},X_{j}\rangle=\sum_{j=1}^{d}\,\langle[[X_{i},\widebar{X}_{i}]_{\mathfrak{m}},\widebar{X}_{j}]_{\mathfrak{m}},X_{j}\rangle,$
while $\operatorname{ad}_{\mathfrak{k}}$-invariance implies that
$\langle X_{i},[[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{k}},X_{j}]_{\mathfrak{m}}\rangle=\langle[X_{i},[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{k}}]_{\mathfrak{m}},X_{j}\rangle.$
It follows that the last $j$-sum vanishes, and as $B(X_{i},\widebar{X}_{i})=(-1)^{[i]}B(\widebar{X}_{i},X_{i})$, the formula in Theorem 3.13 follows. ∎
@@ 3.5 Naturally reductive metrics
We say that the $G$-invariant metric $g$ on the homogeneous superspace $M=G/K$ is naturally reductive if
@@
for all $X,Y,Z\in\mathfrak{m}$. This mimics the definition of a naturally reductive metric in classical homogeneous geometry [DZ79]. Such metrics have been used extensively in the study of the Einstein equation, spectral geometry, the prescribed Ricci curvature problem, and other topics; see, e.g., [DZ79, GS10, Lau20, AGP20]. When $g$ is naturally reductive, the curvature formulas obtained above simplify substantially.
@@
Let $g$ be naturally reductive. Then, $U=0$.
@@
Since this holds for all $Z$, the result follows from the non-degeneracy of $\langle-,$-$\rangle$. ∎
@@
Let $M=G/K$ be a homogeneous superspace with $G$-invariant graded naturally reductive Riemannian metric $g$, where $G=(G_{0},\mathfrak{g})$ is a connected Lie supergroup, $\mathfrak{g}$ a compact real form of a basic classical Lie superalgebra, and $K$ a connected closed Lie subsupergroup of $G$. Let $\{X_{1},\ldots,X_{d}\}$ be a $g$-normalised homogeneous basis for $\mathfrak{m}=T_{K}(M)$, with right dual basis $\{\widebar{X}_{1},\ldots,\widebar{X}_{d}\}$. Then, for every $i,j,k,\ell\in\{1,\ldots,d\}$,
@@ 3.6 Diagonal metrics
Recall that $\mathfrak{g}$ admits the $Q$-orthogonal decomposition $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{m}$. As $K$ is connected, this means that $\mathfrak{m}$ is an $\operatorname{ad}_{\mathfrak{k}}$-representation, that is, $\mathfrak{k},\mathfrak{m}]\subseteq\mathfrak{m}$. In the following, we will assume $\mathfrak{m}$ decomposes into $Q$-orthogonal $\operatorname{ad}_{\mathfrak{k}}$-irreducible representations,
@@
As we will discuss in the following, the curvature formulas obtained in Section 3.3 simplify quite dramatically for diagonal metrics. We will frequently use an ordered basis for $\mathfrak{m}$ that is adapted to the decomposition (3.24). For each $i=1,\ldots,s$, we thus let $\mathfrak{M}_{i}=\{e^{i}_{\alpha}\,|\,\alpha=1,\ldots,\dim(\mathfrak{m}_{i})\}$ be an ordered $Q$-normalised homogeneous basis for $\mathfrak{m}_{i}$ such that the corresponding matrix representation of $Q$ is of the form (3.11), and let $\{\bar{e}_{\alpha}^{i}\,|\,\alpha=1,\ldots,\dim(\mathfrak{m}_{i})\}$ be the right $Q$-dual basis, defined by
@@
Let $g$ be a $G$-invariant graded Riemannian metric on $M$. Then, by Theorem 3.6, $g$ is identified with an $\operatorname{Ad}_{K}$-invariant scalar superproduct $\langle-,$-$\rangle$ on $\mathfrak{m}$. Since $K$ is connected, the $\operatorname{Ad}_{K}$-invariance is equivalent to the $\operatorname{ad}_{\mathfrak{k}}$-invariance of $\langle-,$-$\rangle$ at the level of Lie superalgebras. By Schur’s lemma, the metric $g$ is therefore of the form
$\langle-,-\rangle=\sum_{i=1}^{s}x_{i}Q|_{\mathfrak{m}_{i}},\qquad x_{1},\ldots,x_{s}\in\mathbb{R}^{\times},$ (3.26)
@@ 3.6.1 Structure constants
Using $\bar{e}^{j}_{\beta},\bar{e}_{\alpha}^{i}]_{\mathfrak{m}_{k}}=\sum_{\gamma\in I^{k}}Q([\bar{e}^{j}_{\beta},\bar{e}_{\alpha}^{i}]_{\mathfrak{m}},\bar{e}_{\gamma}^{k})e_{\gamma}^{k}$, we see that
@@
It remains to show that $[ijk]$ is independent of the choice of basis for $\mathfrak{m}_{i}$. Suppose there is another $Q$-normalised homogeneous basis $\{f_{\beta}^{i}\,|\,\beta\in I^{i}\}$ for $\mathfrak{m}_{i}$, with right dual basis $\{\bar{f}_{\beta}^{i}\,|\,\beta\in I^{i}\}$, and let $A^{i}$ and $B^{i}$ be the (even) transition matrices such that
@@
Since $\mathfrak{g}$ admits the $Q$-orthogonal decomposition $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{m}$ and $\mathfrak{k},\mathfrak{m}]\subseteq\mathfrak{m}$, we have
@@ 3.6.2 Ricci curvature
For diagonal metrics, we now derive expressions for the Ricci coefficients $r_{i}$, $i=1,\ldots,s$, in (3.27). Although the superdimensions $d_{i}$ may be 0 (unlike the corresponding ordinary dimensions), the formula for the Ricci coefficients in Theorem 3.26 below is analogous to the corresponding classical result; cf. [PS97, Lemma 1.1] and [PR19, Lemma 3.3]. If $\dim(\mathfrak{m}_{\bar{1}})=0$, then Theorem 3.26 gives a Ricci curvature formula for a pseudo-Riemannian manifold.
@@
For each $i=1,\ldots,s$, let $\{X^{i}_{\alpha}\,|\,\alpha\in I^{i}\}$ be a $g$-normalised basis for $\mathfrak{m}_{i}$, with right $g$-dual basis $\{\widebar{X}_{\alpha}^{i}\,|\,\alpha\in I^{i}\}$, and let $\{X_{1},\dots,X_{d}\}$ be a corresponding basis for $\mathfrak{m}$ adapted to the decomposition (3.24). Then, $Q$-orthogonality implies that $\langle X^{i}_{\alpha},X^{j}_{\beta}\rangle=0$ for $i\neq j$ and any $\alpha,\beta$. For each $i=1,\ldots,s$, we also introduce the renormalised basis $\{e^{i}_{\alpha}=\sqrt{\tilde{x}_{i}}X_{\alpha}^{i}\,|\,\alpha\in I^{i}\}$ for $\mathfrak{m}_{i}$, whose right $Q$-dual basis is seen to be given by $\{\bar{e}_{\alpha}^{i}=\epsilon_{i}\sqrt{\tilde{x}_{i}}\widebar{X}^{i}_{\alpha}\,|\,\alpha\in I^{i}\}$.
@@
where the second, second-to-last and last equalities follow from $X_{\alpha}^{i},X_{j}]_{\mathfrak{m}}=\sum_{k=1}^{d}\,\langle X_{k},[X_{\alpha}^{i},X_{j}]_{\mathfrak{m}}\rangle\,\widebar{X}_{k}$, $\bar{e}_{\beta}^{j},\bar{e}^{i}_{\alpha}]_{\mathfrak{m}_{k}}=\sum_{\gamma\in I^{k}}Q\big([\bar{e}_{\beta}^{j},\bar{e}^{i}_{\alpha}]_{\mathfrak{m}_{k}},\bar{e}_{\gamma}^{k}\big)e_{\gamma}^{k}$ and (3.6.1), respectively. Taking the graded sum over the $Q$-normalised homogeneous basis for $\mathfrak{m}_{i}$, we obtain
@@
In the homogeneous $\mathfrak{m}_{i}$-basis $\{e^{i}_{\alpha}\,|\,\alpha\in I^{i}\}$ introduced in the proof of Theorem 3.26, there exists at least one even vector, because $d_{i}=0$. Let us select one such vector and denote it by $e_{\alpha}^{i}$. Since (3.35) holds, we have
@@ 4 Flag supermanifolds
In this section, we introduce a class of homogeneous superspaces called flag supermanifolds. In analogy with the construction of generalised flag manifolds in terms of so-called painted Dynkin diagrams in the non-super case [Arv03], we introduce what we call circled Dynkin diagrams to describe the flag supermanifolds in Section 4.1. In Section 4.2, we identify the compact real forms of the Lie superalgebras of types $A$ and $C$. In Section 4.3, Section 4.4 and Section 4.5, we then classify Einstein metrics on flag supermanifolds constructed from circled Dynkin diagrams of types $A$ and $C$. We refer to [Kac77] for Dynkin diagrams of complex Lie superalgebras, to [Ser83, Par90, Chu13] for classifications of real forms of Lie superalgebras, and to [Kac78] for the representation theory of complex Lie superalgebras.
@@ 4.1 Circled Dynkin diagrams
$\mathfrak{g}_{\alpha}^{\mathbb{C}}:=\big\{X\in\mathfrak{g}^{\mathbb{C}}\,|\,[H,X]=\alpha(H)X,\ \text{for all}\ H\in\mathfrak{h}^{\mathbb{C}}\big\}.$
@@
$\mathfrak{m}=\operatorname{span}_{\mathbb{R}}\big\{X\in\mathfrak{m}_{\bar{0}}^{\mathbb{C}}\,|\,X^{*}=-X\big\}\oplus\operatorname{span}_{\mathbb{R}}\big\{\sqrt{\imath}X\in\mathfrak{m}_{\bar{1}}^{\mathbb{C}}\,|\,X^{*}=-X\big\},$ (4.3)
@@
$\{A_{\alpha},B_{\alpha}\,|\,\alpha\in\Delta_{M}\cap\Delta_{\bar{0}}^{+}\}\cup\{\sqrt{\imath}A_{\alpha},\sqrt{\imath}B_{\alpha}\,|\,\alpha\in\Delta_{M}\cap\Delta_{\bar{1}}^{+}\}.$
In the non-super case, the above construction is encoded in a so-called painted Dynkin diagram, where the nodes of $\Pi_{M}$ are painted black while the other nodes of $\Pi$ remain white. In the super-setting, however, black nodes typically represent non-isotropic odd simple roots, so to avoid confusion, we shall circle the nodes of $\Pi_{M}$, thereby obtaining a circled Dynkin diagram of $G$. In addition, a node marked with a cross, $\otimes$, denotes an isotropic odd simple root, in accordance with the standard Dynkin diagram conventions for basic classical Lie superalgebras. The uncircled nodes of $\Pi$ form a simple root system of $\mathfrak{k}_{\mathrm{ss}}^{\mathbb{C}}$.